OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)?

OpenStudy (anonymous):

what's the slope of y = 2x + 3 ?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

just 2 actually. :) and parallel lines have the same slope so the slope of the parallel line is also 2. You can write the equation in point slope form \[y-y_1=m(x-x_1)\] m is the slope, so that's 2 (x_1,y_1) is the point so that's (2, -1)

OpenStudy (anonymous):

so the nwser is 2x+y=-1

OpenStudy (anonymous):

y + 1 = 2(x - 2) y + 1 = 2x - 4 y = 2x - 5 2x - y = 5

OpenStudy (anonymous):

y = 2x – 5 can it be written like this

OpenStudy (anonymous):

yes that's slope intercept form. y + 1 = 2(x - 2) is point slope form 2x - y = 5 is standard form 3 different ways to write the same equation

OpenStudy (anonymous):

thanks very much can you help me with some more